OpenStudy (anonymous):

A future of bacteria originally numbers 500. After 2 hours, there are 1500 bacteria in the culture. Assuming exponential growth, how many are there after 6 hours? ***Not sure how to solve this.... How would you start? :/ Please explain! Thanks a bunch! :)

OpenStudy (mertsj):

What is the formula for exponential growth?

OpenStudy (anonymous):

is it P=Poa^t ?

OpenStudy (anonymous):

oh oops.. sorry P=Po(2)^t ?

OpenStudy (anonymous):

but that's for doubling right? :/

OpenStudy (mertsj):

Yes. Use the first equation and find a

OpenStudy (mertsj):

1500=500a^2

OpenStudy (anonymous):

okay 3=a^2 so a= √3 ?

OpenStudy (mertsj):

yes. Now replace t with 6 and find P

OpenStudy (mertsj):

P=500(sqrt3)^6

OpenStudy (anonymous):

P=500(√3)^6=4.22 ?

OpenStudy (anonymous):

or did i do it wrong? do I do the √3^6 first? if so, that gets 27 x 500= 13500?? not sure which way? because the 4.22 looked a bit off haha

OpenStudy (mertsj):

powers first

OpenStudy (anonymous):

okay so P=500(√3)^6 =500(27)=13,500 so there will be 13,500 bacteria in the culture after 6 hours? is that right? :)